OpenStudy (anonymous):

sqrt 2x-1 + sqrt x-4 = 4

OpenStudy (anonymous):

\[\sqrt{2x-1}+\sqrt{x-4}=4\]

OpenStudy (anonymous):

We want to use the difference of two squares: (a+b)(a-b)=a^2-b^2. Multiply both sides by \[(\sqrt{2x-1}-\sqrt{x-4})\]And we get: \[(\sqrt{2x-1}+\sqrt{x-4})(\sqrt{2x-1}-\sqrt{x-4})=4(\sqrt{2x-1}-\sqrt{x-4})\] \[(2x-1)-(x-4)=4\sqrt{2x-1}-4\sqrt{x-4} \rightarrow x+3=4\sqrt{2x-1}-4\sqrt{x-4}\] Now we have two equations which we can work with:\[\sqrt{2x-1}+\sqrt{x-4}=4\]\[4\sqrt{2x-1}-4\sqrt{x-4}=x+3\] Add four times the first equation to the second equation and we get:\[8 \sqrt{2x-1}=x+19\]\[64(2x-1)=x^2+38x+361\]\[x^2-90x+425=0\] This gives x=5, x=85.